OpenStudy (anonymous):

Let f(x)=(2x^2-x). Give f '(1).

OpenStudy (anonymous):

I'm a little stumped on this "Give f '(1) problems.

Parth (parthkohli):

First, find the derivative. What is \(f'(x)\)?

OpenStudy (anonymous):

1?

Parth (parthkohli):

No, no. It'd be\[f'(x) = {d \over dx}2x^2 - x\]Do you know what derivative is?

OpenStudy (anonymous):

Let f(x)=(2x^2-x). Give f '(1). f'(x) = 4x f'(1) = 4(1) = 4 . am i right?

OpenStudy (anonymous):

Well we are just starting the chapter on derivatives, so I'm still figuring them out. Mostly we've been dealing with finding the slope of tangent and normal lines.

Parth (parthkohli):

@febylailani I should point out that,\[f'(x) = 4x - 1\]

OpenStudy (anonymous):

3

Parth (parthkohli):

@cuzzin That includes differentiation too!

Parth (parthkohli):

\[f'(x) = 4x - 1\]\[f'(1) = 4(1) - 1\]

OpenStudy (anonymous):

I think you should write it as: \[\frac{d}{dx} \left(2x^{2}-x\right) or \frac{d}{dx} 2x^2-\frac{d}{dx} x\] Or people would get confused,which was evident :)

OpenStudy (anonymous):

why to make things more complicated

Parth (parthkohli):

@Omniscience is pretty well till now, and then the power rule!:)

OpenStudy (anonymous):

ah!! yup! sorry i forgot. careless -_- aaaaaccckkkkk

OpenStudy (anonymous):

Ok, where does the (4x-1) come from?

OpenStudy (anonymous):

from differential.

OpenStudy (anonymous):

but differential is the short way. before that, there's limit. the formula is a.n.x^(n-1). i.e. f(x) = 2x^3 ; a=2 ; n = 3 so, f'(x) = 2.3.x^(3-1) = 6x^2 cmiiw